OpenStudy (anonymous):

Solve: -4(x + 10) - 6 = -3(x - 2)

OpenStudy (anonymous):

Multiply the constants out to find \[-4x-40-6=-3x+6\] Then collect like terms to find \[-4x+3x=40+6+6\] Leave it to you from there!

OpenStudy (anonymous):

OMG! Thank you so much I am really bad at algebra your a life saver!